Q:-A wire of length 'l' having tension T and radius 'r' vibrats with fundamental frequency 'f'.Another wire of the same metal with length '2l' having tension 2T and radius 2r will vibrate with fundamental frequency:


A.f


B.2f


C.f/2√2


D.(f/2)√2

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago
Dear Tapasranjan let the density of the material is ? so mass per unit length (µ) =??r^2 f=1/2L v(T/µ) =1/2L v(T/??r^2) now mass per unit length for second wire =??(2r)^2 =4??r^2 f' =1/4L v(2T/4??r^2) = f/2v2 Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE & AIEEE preparation. All the best. Regards, Askiitians Experts Badiuddin